CSE 663, Advanced Knowledge Representation, Spring 2002 Professor Stuart C. Shapiro Homework 3

Due at Start of Class, September 26, 2002

1. (a) Represent the following^{1} in SNePS 2.6. i. Mammals have hair and give milk. ii. Birds have feathers, fly, and lay eggs. iii. Carnivores are mammals, eat meat, have pointed teeth, have claws, and their eyes point

forward. iv. Ungulates are mammals, have hoofs, and chew cud. v. Cheetahs are carnivores, are tawny in color, and have dark spots. vi. Tigers are carnivores, are tawny in color, and have black stripes. vii. Giraffes are ungulates, have long legs, have long necks, are tawny in color, and have dark

spots. viii. Zebras are carnivores, are white, and have black stripes. ix. Ostriches are birds, do not fly, have long legs, have long necks, and are black and white. x. Penguins are birds, do not fly, swim, and are black and white. xi. Albatrosses are birds and are good fliers.

(b) Using path-based reasoning, and wire-based reasoning if necessary, write: i. a find expression that finds the animal which has long legs and feathers. ii. a find expression that finds the animal which is tawny in color, has dark spots, chews cud

and gives milk.^{1}

2. Do Problem 1 in SNePS 3.

What to hand in: Prepare two demo files, one for SNePS 2.6 and one for SNePS 3, that build the knowledge base and perform the finds. Hand in a listing produced by running demo on these two files, and email the two files to Prof. Shapiro. Make the subject of the two emails “CSE 663 HW3 Q1” and “CSE 663 HW3 Q2”.

1Based on P. H. Winston, Artificial Intelligence, Second Edition, Addison-Wesley, Reading, MA, 1984, pp. 182–185.